OpenStudy (imnotgoodatmath):

What is the value of 10^-3 in standard form?

OpenStudy (imnotgoodatmath):

@zaibali.qasmi

OpenStudy (solomonzelman):

There is a rule, \(\normalsize\color{black}{ a^{-b}=\frac{\Large 1}{\Large a^b} }\). And this \(\normalsize\color{black}{ 10^{-3} }\) is one of the cases where we apply this rule. \(\normalsize\color{blue}{ 10^{-3}=\frac{\Large 1}{\Large 10^3} =\frac{\Large 1}{\Large 1000} =0.001 }\)

OpenStudy (anonymous):

10^-3=\[10^-3=\frac{ 1 }{ 10^3 }=\frac{ 1 }{ 1000 }=0.001\]